### v1

@@ 3.4 Optimising Cavity Length and Corresponding Acceleration Sensitivity for a Particular Momentum Width
We also find in Figure 7 that the maximum optimal Fisher information decreases as the momentum width increases. This is because, as shown in Ref. manju_atomic_2020 , resonant transmission is suppressed and transmission peaks broaden as momentum width $\Delta k$ increases. A decrease in Fisher information corresponds to a worse sensitivity (i.e. a larger $\delta a$), as illustrated in Figure 8 by plotting $\delta a$ as a function of momentum width $\Delta k$.
@@
3.5 Comparing Sensitivities of a MZ Interferometer and an Atomic FPI
@@ 4 Conclusions
This work has focused on the atomic FPI as a space-based accelerometer. When applied as an earth-based accelerometer, the atomic cloud gains significantly more energy in the accelerating period before the first barrier. The barrier height must then be increased to exceed the atomic energy (a necessary condition for the atomic FPI to function), which necessitates a reduction in barrier width and cavity length as per the parameters determined in Ref. manju_atomic_2020 . The barrier width required now is three orders of magnitude smaller than achievable by existing laser focusing methods. To overcome this difficulty, we propose an alternative scheme: initially position the atoms on the right hand side of the barrier in Figure 1, and provide a momentum kick against the direction of the acceleration (towards the barriers). The acceleration will therefore reduce the energy of the atomic cloud, so that an experimentally feasible barrier width may be used.
[…]
Acknowledgments SAH acknowledges support through an Australian Research Council Future Fellowship Grant No. FT210100809. SSS was supported by an Australian Research Council Discovery Early Career Researcher Award (DECRA), Project No. DE200100495. This research was undertaken with the assistance of resources and services from the National Computational Infrastructure (NCI), which is supported by the Australian Government.

### v2

@@ III.4 Optimising Cavity Length and Corresponding Acceleration Sensitivity for a Particular Momentum Width
We also find in Figure 7 that the maximum optimal Fisher information decreases as the momentum width increases. This is because, as shown in Ref. Manju et al. (2020), resonant transmission is suppressed and transmission peaks broaden as momentum width $\Delta k$ increases. A decrease in Fisher information corresponds to a worse sensitivity (i.e. a larger $\delta a$), as illustrated in Figure 8 by plotting $\delta a$ as a function of momentum width $\Delta k$.
@@
III.5 Comparing Sensitivities of a MZ Interferometer and an Atomic FPI
@@ IV Conclusions
This work has focused on the atomic FPI as a space-based accelerometer. When applied as an earth-based accelerometer, the atomic cloud gains significantly more energy in the accelerating period before the first barrier. The barrier height must then be increased to exceed the atomic energy (a necessary condition for the atomic FPI to function), which necessitates a reduction in barrier width and cavity length as per the parameters determined in Ref. Manju et al. (2020). The barrier width required now is three orders of magnitude smaller than achievable by existing laser focusing methods. To overcome this difficulty, we propose an alternative scheme: initially position the atoms on the right hand side of the barrier in Figure 1, and provide a momentum kick against the direction of the acceleration (towards the barriers). The acceleration will therefore reduce the energy of the atomic cloud, so that an experimentally feasible barrier width may be used.
Acknowledgements.
SAH acknowledges support through an Australian Research Council Future Fellowship Grant No. FT210100809. SSS was supported by an Australian Research Council Discovery Early Career Researcher Award (DECRA), Project No. DE200100495. This research was undertaken with the assistance of resources and services from the National Computational Infrastructure (NCI), which is supported by the Australian Government.